### v1

# RAG-Reward: Optimizing RAG with Reward Modeling and RLHF

Abstract
Retrieval-augmented generation (RAG) enhances Large Language Models (LLMs) with relevant and up-to-date knowledge, improving their ability to answer knowledge-intensive questions. It has been shown to enhance both generation quality and trustworthiness. While numerous works have focused on improving retrieval, generation, and evaluation, the role of reward models in reinforcement learning for optimizing RAG and establishing automated benchmarking pipelines remains underexplored. In this paper, we introduce RAG-Reward, a dataset designed to enable hallucination-free, comprehensive, reliable, and efficient RAG. We define four key metrics for assessing generation quality and develop an automated annotation pipeline that leverages multiple LLMs to generate outputs across diverse RAG scenarios. GPT-4o is used to evaluate and construct preference data. Using RAG-Reward, we train reward models and apply reinforcement learning with human feedback (RLHF) to improve LLMs’ effectiveness in RAG. Experimental results show that our reward model achieves state-of-the-art performance on a held-out test set, demonstrating both the effectiveness of our approach and the quality of our dataset. Furthermore, the improved generation quality of the trained policy model highlights the feasibility of using RLHF to enhance RAG pipelines [^1] Our dataset is available at: https://huggingface.co/datasets/HanningZhang/RAG-Reward-Modeling.
RAG-Reward: Optimizing RAG with Reward Modeling and RLHF
Hanning Zhang1, Juntong Song2, Juno Zhu2, Yuanhao Wu2, Tong Zhang1, Cheng Niu2 1University of Illinois Urbana-Champaign, 2NewsBreak {hanning5, tozhang}@illinois.edu {juntong.song, juno, yuanhao.wu, cheng.niu}@newsbreak.com
1 Introduction
Figure 1: An overview of our data labeling method and our experiments based on the preference data in RAG Scenario. We use GPT-4o as the judge to evaluate the quality of the generation from multiple models. We then train the reward models and use them for Reinforcement Learning.
Retrieval-augmented generation (RAG) has been widely adopted in research and real-world applications on domain-specific or knowledge-intensive tasks (Gao et al., 2024; Fan et al., 2024; Yu et al., 2023). By leveraging up-to-date external knowledge, Large Language Models (LLMs) can incorporate relevant information during text generation, significantly mitigating hallucination issues (Zhang et al., 2023b; Peng et al., 2023) and ensuring higher answer quality (Lewis et al., 2020).
However, many open-source LLMs are not yet fully optimized for use in RAG scenarios. Simply applying RAG to these models often leads to suboptimal outcomes. For example, even with access to external knowledge, these LLMs may still generate misinformation (Niu et al., 2024), which can be especially problematic given the timeliness of the generated content. While many LLMs excel in conversational tasks, they frequently struggle in information-intensive scenarios, where both efficiency and simplicity are paramount. Therefore, a solution that can provide effective supervision and evaluation of RAG systems is needed.
Fortunately, significant progress has been made in evaluating the quality of LLM-generated content these days. Unlike many previous benchmark data sets that are created to assess the overall quality of generations (Yang et al., 2018; Kwiatkowski et al., 2019), recent studies suggest that the construction of evaluation data sets tailored to specific domains or scenarios with detailed criteria (Zhu et al., 2024; Friel et al., 2024) can lead to improved in-domain performance. We argue that it is also feasible to build such a preference dataset that reflects the quality of RAG based on multiple key metrics carefully selected and defined by human experts.
At the same time, reward modeling has recently emerged as a widely adopted strategy to align text generation with human values by learning preference signals from human-annotated high-quality data (Ouyang et al., 2022; Lambert et al., 2024). Multiple well-crafted datasets have demonstrated the value of reward models and high-quality preference datasets. For example, HH-RLHF (Bai et al., 2022a) evaluates the helpfulness and harmlessness of the language, PRM800K (Lightman et al., 2023) assesses the stepwise correctness during mathematical reasoning, and RAG-Truth (Niu et al., 2024) detects hallucination occurrence in RAG systems.
While human annotation has proven to be an effective method to evaluate the quality of generated responses and constructing datasets, it is both expensive and time-consuming. These limitations have led many researchers to explore reliable annotation strategies that leverage the capabilities of LLMs. Previous studies have demonstrated the feasibility of using LLMs to assess response quality and build robust reward models (Tan et al., 2024). For example, Bai et al. (2022b) introduced Constitution AI, which uses LLMs to guide the generation of fine-tuning data and provide preference pairs for reinforcement learning. Moreover, LLMs can also be guided by human-defined criteria and examples to provide preference scores tailored to more specific requirements (Sun et al., 2023, 2024).
The use of LLMs for annotation has further proven effective when applied to a broader range of tasks and larger datasets. For example, Cui et al. (2024) introduced Ultra-Feedback, a framework that leverages GPT-4 to assign scores based on attributes such as helpfulness, truthfulness, honesty, and instruction-following ability. LLMs have also shown their effectiveness in providing feedback in a variety of reasoning tasks, including coding and solving mathematical problems (Weyssow et al., 2024; Yuan et al., 2024).
Building on the inspiration from these works, we further harness the capabilities of LLMs to construct a preference dataset tailored to specific RAG frameworks across diverse domains. To achieve this, we develop a novel automatic LLM annotation pipeline, culminating in a high-quality preference dataset, RAG-Reward, explicitly designed for the general RAG task. Figure 1 shows an overview of our framework. Several works are related to this, such as measuring the relevance of query and passage using reward models (Nguyen et al., 2024), and improve trustworthiness of RAG using DPO (Song et al., 2024). However, to the best of our knowledge, we are the first to integrate RAG with the standard reward modeling and RLHF pipeline.
Specifically, we select datasets from domains of Question-Answering, Data-to-Text, and Summarization. First, we sample diverse responses from a pool of 12 open-source and proprietary LLMs, including GPT-4 and the Llama-3 series. For each prompt in the datasets, we randomly select 2 LLMs to generate responses. Next, we use GPT-4o as the judge to compare them based on four key metrics carefully selected by the human experts: hallucination, comprehensiveness, verbosity, and attribution. This enables us to construct preference pairs, consisting of a chosen response and a rejected one, based on majority voting across 4 metrics. Overall, we collect 35K high-quality training samples for the reward model. Evaluations demonstrate that our reward model achieves over 80% accuracy on the held-out test set. Additionally, we develop a policy model using the RAFT algorithm (Dong et al., 2023), leading to notable performance improvements. Our key contributions are summarized as follows:
  • $\bullet$
    We introduce a reward modeling method for the RAG scenario to assess generation quality. Additionally, we release a high-quality dataset of 35K preference annotations to support future research.
  • $\bullet$
    We define a comprehensive set of metrics that effectively evaluate RAG quality and guide the dataset construction process.
  • $\bullet$
    We conducted extensive experiments to evaluate our reward model, train a policy model, and demonstrate the effectiveness of our approach in enhancing RAG performance.
2 Related Work
2.1 Reward Modeling for Alignment and Reinforcement Learning
Training reward models have become a widely used approach to align language models with human preference (Ouyang et al., 2022). The alignment can enhance various aspects of LLM performance, such as increasing their trustworthiness and helpfulness (Bai et al., 2022a; Wang et al., 2023; Cui et al., 2024), or improving their problem-solving abilities (Dai et al., 2024; Yuan et al., 2024; Zhang et al., 2024a). The reward signal can be trained as a discriminative model to generate a scalar value (Bradley and Terry, 1952), or directly generated as critics from language models (Zhang et al., 2024b; Zheng et al., 2023). Many high-quality datasets for reward modeling have been introduced, such as HH-RLHF (Bai et al., 2022a), Ultra-Feedback (Cui et al., 2024), Code-UltraFeedback (Weyssow et al., 2024), Ultra-Interact (Yuan et al., 2024), and PKU-SafeRLHF (Ji et al., 2024), which could be either labeled by human or by powerful LLMs.
Reinforcement Learning from human feedback (RLHF) is a widely used strategy to enhance policy models after the reward model is developed (Kaufmann et al., 2024). RLHF plays a critical role in aligning LLMs with human values and achieving improved performance (Christiano et al., 2023). Proximal Policy Optimization (PPO) is a commonly used algorithm for alignment tasks to enhance the policy models (Schulman et al., 2017), although it is computationally intensive. Consequently, several alternative, more efficient algorithms have been proposed, such as Direct Preference Optimization (DPO) (Rafailov et al., 2024), Kahneman-Tversky Optimization (KTO) (Ethayarajh et al., 2024), and Rejection Sampling Fine-tuning (RAFT) (Dong et al., 2023). These approaches have also been widely adopted in state-of-the-art models like Llama-3 (Grattafiori et al., 2024) and Qwen-2 (Yang et al., 2024a).
Concurrently, the work introduced in Jin et al. (2024) utilizes existing reward models to evaluate Question-Answering tasks in RAG scenarios with fine-grained metrics, highlighting the limitation of the general reward models. It also shows the feasibility of constructing RAG scenario data using Large Language Models. Our work is built upon these existing works to train the RAG-specific reward model and use it for alignment training.
2.2 Large Language Models and Retrieval Augmented Generation
Retrieval-Augmented Generation (RAG) has proven to be an effective method for enhancing language models with real-world knowledge to address a wide range of tasks, thereby improving the accuracy and credibility of the generated output (Lewis et al., 2021). In the era of LLMs, which possess a strong ability to understand and utilize in-context information, RAG can significantly enhance their capabilities (Fan et al., 2024; Gao et al., 2024). RAG addresses common challenges of LLMs, such as hallucinations and outdated knowledge, by grounding their outputs in external knowledge bases (Peng et al., 2023; Li et al., 2024). RAG-based LLMs can be trained to effectively adapt and integrate retrieved information (Schick et al., 2023; Shao et al., 2023), or use training-free methods that directly insert the retrieved context into the prompt (Ram et al., 2023). These LLMs have been widely adopted in real-world applications. For instance, retrievers are integrated into LLM-based chatbots to increase the helpfulness and trustworthiness of the conversations (Komeili et al., 2021). RAG-based models have also been deployed as domain-specific experts, such as finance (Zhang et al., 2023a) and medicine (Xiong et al., 2024).
In this project, we are the first to systematically construct RAG-scenario datasets and develop reward models, paving the way for evaluating and enhancing the generation quality of LLMs within the RAG framework.
Figure 2: An illustration of our data annotation method. Given a sample and two responses, we prompt GPT-4o to provide a judgment based on each metric separately. We then aggregate the results and construct the pairs.
3 Dataset Construction
We construct our dataset based on existing RAG datasets to ensure its relevance and applicability. To reflect the diverse use cases of RAG scenarios, we include three common types: Question Answering, Data-to-Text and Summarization. Specifically, we use WebGLM (Liu et al., 2023), Yelp (Yelp, 2021), and XSum (Narayan et al., 2018) as experimental datasets, each dataset corresponding to one of the three RAG scenarios. Examples of these data sets are presented in Table 2.
Figure 3: The percentage contribution of different metrics for the chosen responses. Each bar shows the percentage of the chosen responses are preferred on that metric.
For the WebGLM dataset, LLMs are tasked with reasoning over web-retrieved reference data to answer real-world questions, generating concise responses in a few sentences. For the Yelp dataset, our experiments focus on data from the restaurant category, represented in JSON format. Each sample includes information such as a restaurant’s location and ambiance. Based on the structured JSON input, LLMs generate descriptive text about the restaurant. The XSum dataset contains diverse articles from the British Broadcasting Corporation (BBC), with models tasked with summarizing these articles. These three datasets cover a broad range of circumstances, ensuring that the reward model trained on them can significantly improve the development and evaluation of RAG systems. Table 1 presents the number of data samples used in our experiments.
When evaluating the quality of the responses, we consider the following metrics:
Hallucination:
The models should generate responses strictly based on the context provided, without introducing information not grounded in the retrieval context. If the retrieval context contradicts the model’s parametric knowledge, the model should adhere to the retrieval reference, ensuring that the response is accurate and contextually relevant.
Comprehensiveness:
The response should fully utilize the context provided by the retrieved content and address all aspects of the instruction. This requires the model to extract and integrate all relevant information from the retrieval context to ensure the response is thorough and complete.
Verbosity:
While the response should be detailed and comprehensive, it should also be concise and straight to the point. Striking the right balance between detail and brevity is essential for delivering informative responses without overwhelming the user.
Attribution:
This metric is specifically applied to the WebGLM-QA dataset to ensure the generations are trustworthy and verifiable. The response should explicitly refer to the context retrieved to improve credibility and allow users to trace information sources.
Dataset Number of Samples
Training Dev Testing
WebGLM 11000 1000 500
Yelp 12000 1000 500
XSum 12000 1000 500
Table 1: The number of preference pairs that we construct from the 3 datasets in our experiments.
3.1 Dataset Sampling
We utilize a combination of open source instruction models, the GPT-3.5 (Brown et al., 2020) and GPT-4 (OpenAI et al., 2024) series to generate data, ensuring diversity and inclusion of both high-quality and relatively low-quality responses. The open-source models consist of various sizes of the instruction-tuned versions of Llama-3, Llama-3.1, Llama-3.2 (Grattafiori et al., 2024), Llama-2 (Touvron et al., 2023), Qwen-2 (Yang et al., 2024a), InternLM-2 (Cai et al., 2024), and Mistral (Jiang et al., 2023).
In total, we include 12 candidate models for generation. For each question and its corresponding reference in the dataset, we randomly select two models’ generations to form preference pair.
3.2 Dataset Labeling
Dataset Data Example
WebGLM (Liu et al., 2023) Question: Why are different tiers (regular < mid < premium) of gas’ prices almost always 10 cents different? References: [The gap between premium and regular gas has…, According to national averages, the price…] Answer: The 10 cent difference between the different tiers of gas prices is likely due to a convention…
Our Prompt: Answer the following question: {question} Your response should be based on the following passages: {passages} When you respond, you should refer to the source of information…
Yelp (Yelp, 2021) [ Name: The Green Pheasant Address: 215 1st Ave S City: Nashville State: TN Attributes:{ HappyHour: True, DogsAllowed: False, … } ]
Our Prompt: Write an overview about the following business based only on the provided structured data in the JSON format…
XSum (Narayan et al., 2018) Document: The full cost of damage in Newton Stewart, one of the areas worst affected, is still being assessed. Repair work is ongoing in Hawick and many roads… Summary: Clean-up operations are continuing across the Scottish Borders and Dumfries and Galloway after…
Our Prompt: Summarize the following document: {document}…
Table 2: Illustration and statistics of the original datasets and the prompts used to construct the preference data. For WebGLM, LLMs will generate responses based on the reference. For Yelp, LLMs will convert the JSON data into a descriptive overview. For XSum, LLMs will summarize the given document.
We use GPT-4o (OpenAI et al., 2024) to label the data. An illustration of our labeling methods is shown in Figure 2. Given a question and a pair of responses from different models, we prompt GPT-4o to compare and select the preferred response.
Specifically, we ask GPT-4o to compare the responses based on the four metrics outlined earlier, assessing them individually. The responses are then ranked according to these evaluations, with the response exhibiting more preferred metrics being chosen as the selected response, and the other marked as rejected. In cases where both responses receive an equal number of preferred metrics, the data is discarded. This approach acknowledges real-world scenarios where responses are rarely perfect, and trade-offs are often necessary.
3.3 Dataset Composition
We analyze the composition of the constructed dataset by examining how each metric contributes to the formation of the preference data. Since our selection process is based on the majority of preferred metrics, the chosen responses may not excel across all four dimensions. For example, a chosen response could properly refer to the context, cover all the details specified in the question without hallucinations, but be too verbose. To better understand how each metric influences the overall preference decision, we evaluate the quality of each metric in relation to the chosen responses.
We calculate the preferred percentage of each metric, and the results are shown in Figure 3. Firstly, we observe a high preferred percentage of each metric. Overall, the average preferred rate across the four dimensions is nearly 90%, demonstrating the high quality of the chosen response on the 4 metrics with respect to the evaluated metrics. We also observe varying preferred rates between the metrics. Specifically, hallucination and attribution metrics exhibit extremely high preferred percentages, while comprehensiveness and verbosity show a relatively lower rate. This difference suggests that the reward models trained on our preference data are better able to differentiate issues related to hallucinations and attribution in RAG scenarios. In contrast, the reward model might not prioritize comprehensiveness and verbosity metrics in their evaluations.
4 Dataset Evaluation
4.1 Self Evaluation
WebGLM Yelp XSum Avg
Consist. 92.8 92.6 94.8 93.4
Table 3: The consistency rate for the self-evaluation of the constructed dataset. We measure whether the GPT-4o agrees with the comparison results previously made.
Figure 4: An illustration of our self-evaluation method. For a sample of the constructed data pairs, we provide GPT-4o with the previous judgment, and ask it to re-evaluate the comparison result.
In this subsection, we evaluate the consistency of the evaluations provided by GPT-4o, a key metric commonly used to demonstrate the reliability of the annotations. To objectively assess the quality of our constructed dataset, we design a self-evaluation method that measures the consistency of GPT-4o’s responses. An illustration of this method is shown in Figure 4. Specifically, for each prompt, a chosen response, and a rejected response in the dataset, we prompt GPT-4o to revisit its previous comparison result and verify whether it maintains its original judgment. We define the consistency rate as the proportion of samples where the evaluation results remain unchanged across both assessments. In this experiment, we randomly select 1,000 samples for re-evaluation.
The results of this experiment are presented in Table 3. We observe a very high consistency rate across the three tasks and the overall consistency rate exceeds 90%. The results demonstrate that GPT-4 produces stable and consistent labels for most of the data according to well-defined criteria, and reflects the quality of the dataset.
4.2 Human Evaluation
WebGLM Yelp XSum Avg
Consist. 0.68 0.65 0.79 0.71
Table 4: The consistency rate between human evaluation and the GPT-4o labeled dataset.
Models WebGLM Yelp XSum Average
internlm2-7b-reward (Cai et al., 2024) 77.2 83.0 74.0 78.1
UltraRM-13b (Cui et al., 2024) 78.8 82.0 73.2 78.0
FsfairX-LLaMA3-RM-v0.1 (Dong et al., 2024) 74.4 82.8 75.2 77.5
Eurus-RM-7b (Yuan et al., 2024) 75.4 82.6 74.0 77.3
llama-3-tulu-2-8b-uf-mean-rm (Ivison et al., 2024) 75.0 78.2 72.2 75.1
Llama-3-OffsetBias-RM-8B (Park et al., 2024) 68.6 79.0 72.0 73.2
URM-LLaMa-3.1-8B (Lou et al., 2024) 66.4 79.0 67.6 71.0
GRM-Llama3-8B-rewardmodel-ft (Yang et al., 2024b) 66.0 74.0 72.8 70.9
QRM-Llama3.1-8B-v2 (Dorka, 2024) 69.2 72.2 69.2 70.2
Skywork-Reward-Llama-3.1-8B-v0.2 (Liu et al., 2024) 67.4 72.8 67.8 69.3
GRM-Llama3.2-3B-rewardmodel-ft (Yang et al., 2024b) 67.4 68.0 70.6 68.7
Table 5: The evaluation results of the existing reward models on the 3 tasks. They achieve SOTA performance on chatting, safety, and reasoning evaluation, but fall short on RAG tasks.
We also performed human evaluations to assess the alignment of GPT annotations with human preferences. Specifically, we randomly select 100 samples with paired responses from each dataset and ask the annotators to evaluate using the same pipeline described in Section 3.2 and illustrated in Figure 2. Annotators compare the responses based on each metric and determine the preference pair. We calculate the agreement ratio between the human annotators and GPT-4o in the preference pairs, and the results are shown in Table 4. We observe a promising overall agreement rate of 71%, showing the potential of AI-assisted labeling on multiple RAG tasks.
5 Limitation of Existing Reward Models
In this section, we evaluate several existing reward models in our test set. We select models from RewardBench (Lambert et al., 2024) known for their strong performance in assessing aspects such as helpfulness, safety, and reasoning. We examine their performance on diverse RAG scenarios using our curated test set, and demonstrate their limitation of the evaluation on RAG domains.
The experiment results are shown in Table 5. While many of the listed reward models achieve accuracy higher than 90% in evaluating chat, safety, and reasoning tasks, their overall accuracy in RAG scenarios is less than 80%. This underscores a significant gap between mainstream reward models and the unique requirements of RAG tasks. We also observe inconsistent performance across tasks for the reward models, with better results on Data-to-Text tasks (e.g., the Yelp dataset) compared to the other two types, suggesting that current reward models are not uniformly capable across different RAG scenarios. Interestingly, several reward models that achieve state-of-the-art (SOTA) performance in reasoning and safety evaluations, as shown on the leaderboard [^2] https://huggingface.co/spaces/allenai/reward-bench, do not perform well on RAG tasks. For example, models like URM-LLaMa-3.1-8B (Lou et al., 2024) and Skywork-Reward-Llama-3.1-8B-v0.2 (Liu et al., 2024) underperform in the RAG domain.
In contrast, UltraRM-13b (Cui et al., 2024), which performs suboptimally in reasoning and safety evaluations but excels in assessing helpfulness and instruction-following, achieves near-top accuracy on RAG tasks. This suggests that reward models trained primarily on reasoning tasks may not generalize effectively to evaluating RAG-specific generations. Domain-specific training data are therefore essential to address this gap and improve RAG performance evaluation.
6 Experiments
We conduct both reward model training and reinforcement learning using our RAG-Reward dataset. In total, 35K preference pairs are used for reward modeling (see Table 1). Additionally, we create a 3K-sample development set for sampling and learning during RLHF training. To evaluate the performance of the policy and reward models, a held-out test set of 1.5K samples is used.
6.1 Reward Modeling
WebGLM Yelp XSum Average
Acc. 84.8 88.2 78.4 83.8
Table 6: The evaluation results of the reward model on the 3 tasks. The accuracy is calculated as the proportion of test samples where the reward model assigns a higher score to the chosen response than to the rejected response.
We adopt the common approach to train the Bradley-Terry reward model (Bradley and Terry, 1952; Ouyang et al., 2022) to learn the reward signal from the preference data. Specifically, we use Llama-3.1-8B-Instruct (Grattafiori et al., 2024) as the base model for training. We train the reward model with a learning rate of $2e^{-6}$, a global batch size of 64, a max length of 4096, and an epoch of 1 on 4 H100-80G GPUs.
During the test stage, each test sample contains a chosen response and a rejected response. The accuracy is calculated as the proportion of test samples in which the reward model assigns a higher score to the chosen response than to the rejected one. The detailed results are shown in Table 6. We observe a high accuracy of 83.8% for the reward model, demonstrating its effectiveness in aligning with the intended criteria. Compared to the strong reward models in Table 5, our model achieves the highest accuracy, highlighting the advantage of leveraging RAG-specific reward data.
Furthermore, we observe a consistent accuracy across the 3 tasks, indicating that the reward model could jointly learn the preference signal from diverse tasks and domains. Notably, the reward model achieves the highest accuracy on the Data-to-Text task, while its performance is relatively lower on the Summarization task. This difference suggests that comparing structured data with text data is easier for the reward model, while evaluating the quality of a summarization poses a greater challenge. This observation aligns with our intuition and expectations
6.2 Preference Alignment
Mistral-7B-v0.1 Llama-3.2-3B
WebGLM 74.8 65.2
Yelp 66.4 68.6
Xsum 72.8 65.4
Average 71.3 66.4
Table 7: The RLHF results on the 3 tasks. The win rate is calculated as the proportion of test samples where the reward model assigns a higher score to the response generated by the post-trained policy model.
Mistral-7B-v0.1 Llama-3.2-3B
WebGLM 66.8 65.8
Yelp 54.4 57.8
XSum 68.2 72.6
Average 63.1 66.2
Table 8: The RLHF results on the 3 tasks. The win rate is calculated as the proportion of test samples where the GPT-4o prefers the response generated by the post-trained policy model.
We adopt the RAFT algorithm (Dong et al., 2023) to perform the preference alignment. RAFT utilizes the reward model to select the response with the highest reward score from $N$ candidate responses, and then fine-tunes the policy model on this selected set of responses.
We set $N=16$ in our experiments. We use Llama-3.2-3B-Instruct (Grattafiori et al., 2024) and Mistral-7B-Instruct-v0.1 (Jiang et al., 2023) as initial policy models for sampling, followed by RAFT training. Both Mistral and Llama models are fine-tuned with a learning rate of $5e^{-6}$, an epoch of 1, a packing length of 4096, and a global batch size of 16, using the axolotl package [^3] https://github.com/axolotl-ai-cloud/axolotl.
To measure the improvement brought by alignment training, we first sample responses from our held-out test set using both the initial models and the post-trained policy models. Thus, for each prompt, we have paired responses from the two models. These pairs are then evaluated using the reward model, and we calculate the proportion of cases where the responses from the post-trained models are preferred. Furthermore, we ask GPT-4o to compare the pairs based on the criteria introduced in Section 3. For both metrics, a baseline of 50% indicates no improvement in the policy models. The calculated scores are referred to as the win rate against the initial model.
The experiment results are shown in Table 7 and Table 8. We observe a clear improvement in the policy models after a single iteration of RAFT. Both the reward model and GPT-4o agree that generations align more closely with the RAG metrics, as the average win rate is significantly above 50%. These results highlight the effectiveness of our dataset and the reward model.
However, we also observe some imbalance in the reward model’s learning compared to GPT-4o. As shown in the tables, the differences in win rates between the two evaluators are minimal for the WebGLM and XSum test sets. However, the difference is more pronounced for the Yelp dataset. This again suggests that the reward model faces greater challenges in evaluating the quality of structured data, such as JSON-format data.
7 Conclusion
In this paper, we introduce RAG-Reward, a high-quality preference dataset designed for Retrieval-Augmented Generation (RAG). Our dataset is generated through a novel automated AI annotation pipeline, leveraging both open-source and proprietary models to enhance generalization and versatility. To ensure fair and reliable evaluations, we use GPT-4o to assess generation quality based on four key metrics carefully selected by human experts. The dataset spans multiple domains, including Question Answering, Data-to-Text, and Summarization, resulting in a large-scale and diverse benchmark. The experimental results show strong alignment with human evaluations, demonstrating the effectiveness of RAG-Reward in reward modeling and reinforcement learning. These findings highlight the potential of our dataset to advance both the evaluation and generation of RAG systems. To foster further research, we publicly release the dataset to the community.