### v2

# RAG-Reward: Optimizing RAG with Reward Modeling and RLHF

Abstract
Retrieval-augmented generation (RAG) enhances Large Language Models (LLMs) with relevant and up-to-date knowledge, improving their ability to answer knowledge-intensive questions. It has been shown to enhance both generation quality and trustworthiness. While numerous works have focused on improving retrieval, generation, and evaluation, the role of reward models in reinforcement learning for optimizing RAG remains underexplored. In this paper, we introduce RAG-Reward, a framework designed to develop reward models to enable hallucination-free, comprehensive, reliable, and efficient RAG. We define four key metrics to assess generation quality and develop an automated benchmarking pipeline to evaluate the outputs of multiple LLMs across a variety of RAG scenarios. Using RAG-Reward, we train reward models and apply reinforcement learning with human feedback (RLHF) to improve LLMs’ effectiveness in RAG. Experimental results demonstrate that our reward model achieves state-of-the-art performance in automatic benchmarking and aligns closely with human evaluations. Furthermore, the improved generation quality of the trained policy model highlights the feasibility and efficiency of using RLHF to enhance RAG outputs  [^1] Our dataset is available at: https://huggingface.co/datasets/HanningZhang/RAG-Reward-Modeling .
RAG-Reward: Optimizing RAG with Reward Modeling and RLHF
Hanning Zhang1, Juntong Song2, Juno Zhu2, Yuanhao Wu2, Tong Zhang1, Cheng Niu2 1University of Illinois Urbana-Champaign, 2NewsBreak {hanning5, tozhang}@illinois.edu {juntong.song, juno, yuanhao.wu, cheng.niu}@newsbreak.com
1 Introduction
Figure 1: An overview of our data labeling method and our experiments based on the preference data in RAG Scenario. We use o3-mini as the judge to evaluate the quality of the generation from multiple models. We then train the reward models and use them for Reinforcement Learning.
Retrieval-augmented generation (RAG) has been widely adopted in research and real-world applications on domain-specific or knowledge-intensive tasks (Gao et al., 2024; Fan et al., 2024; Yu et al., 2023). By leveraging up-to-date external knowledge, Large Language Models (LLMs) can incorporate relevant information during text generation, significantly mitigating hallucination issues (Zhang et al., 2023b; Peng et al., 2023) and ensuring higher answer quality (Lewis et al., 2020).
However, many open-source LLMs are not yet fully optimized for use in RAG scenarios. Simply applying RAG to these models often leads to suboptimal outcomes. For example, even with access to external knowledge, these LLMs may still generate misinformation (Niu et al., 2024), which can be especially problematic given the timeliness of the generated content. While many LLMs excel in conversational tasks, they often struggle in information-intensive scenarios, where both trustworthiness and efficiency are paramount Wang et al. (2024). Therefore, a solution that can provide effective supervision and evaluation of RAG systems is needed.
Fortunately, significant progress has been made in evaluating the quality of LLM-generated content these days. Unlike many previous benchmark data sets that are created to assess the overall quality of generations (Yang et al., 2018; Kwiatkowski et al., 2019), recent studies suggest that the construction of evaluation data sets tailored to specific domains or scenarios with detailed criteria (Zhu et al., 2024; Friel et al., 2024) can lead to improved in-domain performance. (Friel et al., 2025) shows the feasibility to build such a preference dataset that reflects the RAG quality based on multiple key metrics carefully selected and defined by human experts.
At the same time, reward modeling has recently emerged as a widely adopted strategy to align text generation with human values by learning preference signals from human-annotated high-quality data (Ouyang et al., 2022; Lambert et al., 2024). Multiple well-crafted datasets have demonstrated the value of reward models and high-quality preference datasets. For example, HH-RLHF (Bai et al., 2022a) evaluates the helpfulness and harmlessness of the language, PRM800K (Lightman et al., 2023) assesses the stepwise correctness during mathematical reasoning. And (Song et al., 2024a) develop hallucination detection models using RAG-Truth (Niu et al., 2024) dataset, and achieves hallucination reduction in RAG output.
While human annotation has proven to be effective in evaluating the quality of the generated responses and constructing datasets, it is both expensive and time-consuming. These challenges have led researchers to explore reliable annotation strategies that leverage the capabilities of LLMs. Previous studies have demonstrated the feasibility of using LLMs to assess response quality and build robust reward models (Tan et al., 2024). For instance, Bai et al. (2022b) introduced Constitution AI, which uses LLMs to guide the generation of fine-tuning data and provide preference pairs for reinforcement learning.
Additionally, LLMs can be guided by human-defined criteria and examples to generate preference scores tailored to specific requirements(Sun et al., 2023, 2024). For example, Cui et al. (2024) introduced Ultra-Feedback, a framework that leverages GPT-4 to assign scores based on attributes such as helpfulness, truthfulness, honesty, and instruction-following ability. LLMs have also shown their effectiveness in providing feedback across various reasoning tasks, including coding and solving mathematical problems (Weyssow et al., 2024; Yuan et al., 2024).
Inspired by these works, we developed RAG-Reward, a framework designed to create reward models that enable hallucination-free, comprehensive, reliable, and efficient RAG. The framework leverages the capabilities of LLMs to evaluate RAG responses across diverse domains, train reward models, and enhance RAG response quality through RLHF. Figure 1 provides an overview of our framework.
Several related works have explored various aspects of this approach, such as using reward models to measure the relevance of queries and passages (Nguyen et al., 2024), improving trustworthiness of RAG using DPO (Song et al., 2024b), leveraging LLM to build RAG benchmark dataset (Friel et al., 2025), and comparing efficiency of various LLMs serving as RAG reward models (Jin et al., 2024). However, to the best of our knowledge, this is the first attempt to optimize RAG with reward modeling and RLHF pipeline.
Specifically, we select datasets from domains of Question-Answering, Data-to-Text, and Summarization. First, we sample diverse responses from a pool of 12 open-source and proprietary LLMs, including GPT-4 and the Llama-3 series. For each prompt in the datasets, we randomly select 2 LLMs to generate responses. Next, we use o3-mini as the judge to compare them based on four key metrics carefully selected by the human experts: hallucination, comprehensiveness, verbosity, and attribution. This enables us to construct preference pairs, consisting of a chosen response and a rejected one, based on the weighted selection across 4 metrics. Overall, we collect 35K high-quality training samples for the reward model training. Evaluations demonstrate that our reward model achieves over 80% accuracy on the held-out test set. Additionally, we develop a policy model using the RAFT algorithm (Dong et al., 2023), leading to notable performance improvements. Our key contributions are summarized as follows:
  • $\bullet$
    We introduce a reward modeling method for the RAG scenario to assess generation quality. Additionally, we release a high-quality dataset of 35K preference annotations to support future research.
  • $\bullet$
    We define a comprehensive set of metrics that effectively evaluate RAG quality and guide the dataset construction process.
  • $\bullet$
    We conducted extensive experiments to evaluate our reward model, train a policy model, and demonstrate the effectiveness of our approach in enhancing RAG performance.
2 Related Work
2.1 Reward Modeling for Alignment and Reinforcement Learning
Training reward models have become a widely used approach to align language models with human preference (Ouyang et al., 2022). The alignment can enhance various aspects of LLM performance, such as increasing their trustworthiness and helpfulness (Bai et al., 2022a; Wang et al., 2023; Cui et al., 2024), or improving their problem-solving abilities (Dai et al., 2024; Yuan et al., 2024; Zhang et al., 2024a). The reward signal can be trained as a discriminative model to generate a scalar value (Bradley and Terry, 1952), or directly generated as critics from language models (Zhang et al., 2024b; Zheng et al., 2023). Many high-quality datasets for reward modeling have been introduced, such as HH-RLHF (Bai et al., 2022a), Ultra-Feedback (Cui et al., 2024), Code-UltraFeedback (Weyssow et al., 2024), Ultra-Interact (Yuan et al., 2024), and PKU-SafeRLHF (Ji et al., 2024), which could be either labeled by human or by powerful LLMs.
Reinforcement Learning from human feedback (RLHF) is a widely used strategy to enhance policy models after the reward model is developed (Kaufmann et al., 2024). RLHF plays a critical role in aligning LLMs with human values and achieving improved performance (Christiano et al., 2023). Proximal Policy Optimization (PPO) is a commonly used algorithm for alignment tasks to enhance the policy models (Schulman et al., 2017), although it is computationally intensive. Consequently, several alternative, more efficient algorithms have been proposed, such as Direct Preference Optimization (DPO) (Rafailov et al., 2024), Kahneman-Tversky Optimization (KTO) (Ethayarajh et al., 2024), Group Relative Policy Optimization Shao et al. (2024), and Rejection Sampling Fine-tuning (RAFT) (Dong et al., 2023). These approaches have also been widely adopted in state-of-the-art models like Llama-3 (Grattafiori et al., 2024), and Qwen-2 (Yang et al., 2024a).
Concurrently, the work introduced in Jin et al. (2024) utilizes existing reward models to evaluate Question-Answering tasks in RAG scenarios with fine-grained metrics, highlighting the limitation of the general reward models. It also shows the feasibility of constructing RAG scenario data using Large Language Models. Our work is built upon these existing works to train the RAG-specific reward model and use it for alignment training.
2.2 Large Language Models and Retrieval Augmented Generation
Retrieval-Augmented Generation (RAG) has proven to be an effective method for enhancing language models with real-world knowledge to address a wide range of tasks, thereby improving the accuracy and credibility of the generated output (Lewis et al., 2021). In the era of LLMs, which possess a strong ability to understand and utilize in-context information, RAG can significantly enhance their capabilities (Fan et al., 2024; Gao et al., 2024). RAG addresses common challenges of LLMs, such as hallucinations and outdated knowledge, by grounding their outputs in external knowledge bases (Peng et al., 2023; Li et al., 2024). RAG-based LLMs can be trained to effectively adapt and integrate retrieved information (Schick et al., 2023; Shao et al., 2023), or use training-free methods that directly insert the retrieved context into the prompt (Ram et al., 2023). These LLMs have been widely adopted in real-world applications. For instance, retrievers are integrated into LLM-based chatbots to increase the helpfulness and trustworthiness of the conversations (Komeili et al., 2021). RAG-based models have also been deployed as domain-specific experts, such as finance (Zhang et al., 2023a) and medicine (Xiong et al., 2024).
In this project, we are the first to systematically construct RAG-scenario preference datasets and develop reward models, paving the way for evaluating and enhancing the generation quality of LLMs within the RAG framework.
Figure 2: An illustration of our data annotation method. Given a sample and two responses, we prompt o3-mini to provide a judgment based on each metric separately. We then aggregate the results and construct the pairs.
3 Dataset Construction
We construct our dataset based on existing RAG datasets to ensure its relevance and applicability. To reflect the diverse use cases of RAG scenarios, we include three common types: Question Answering, Data-to-Text and Summarization. Specifically, we use WebGLM (Liu et al., 2023), Yelp (Yelp, 2021), and XSum (Narayan et al., 2018) as experimental datasets, each dataset corresponding to one of the three RAG scenarios.
For the WebGLM dataset, LLMs are tasked with reasoning over web-retrieved reference data to answer real-world questions, generating concise responses in a few sentences. For the Yelp dataset, our experiments focus on data from the restaurant category, represented in JSON format. Each sample includes information such as a restaurant’s location and ambiance. Based on the structured JSON input, LLMs generate descriptive text about the restaurant. The XSum dataset contains diverse articles from the British Broadcasting Corporation (BBC), with models tasked with summarizing these articles. These three datasets cover a broad range of circumstances, ensuring that the reward model trained on them can significantly improve the development and evaluation of RAG systems. Table 1 presents the number of data samples used in our experiments. And examples of these data sets are presented in Table 2.
When evaluating the quality of the responses, we consider the following metrics:
Hallucination:
The models should generate responses strictly based on the context provided, without introducing information not grounded in the retrieval context. If the retrieval context contradicts the model’s parametric knowledge, the model should adhere to the retrieval reference, ensuring that the response is accurate and contextually relevant.
Comprehensiveness:
The response should fully utilize the context provided by the retrieved content and address all aspects of the instruction. This requires the model to extract and integrate all relevant information from the retrieval context to ensure the response is thorough and complete.
Verbosity:
While the response should be detailed and comprehensive, it should also be concise, relevant, and straight to the point. Striking the right balance between detail and brevity is essential to providing informative answers without overwhelming the user.
Attribution:
This metric is specifically applied to the WebGLM-QA dataset to ensure the generations are trustworthy and verifiable. The response should explicitly refer to the context retrieved to improve credibility and allow users to trace information sources.
Dataset Number of Samples
Training Dev Testing
WebGLM 11000 1000 500
Yelp 12000 1000 500
XSum 12000 1000 500
Table 1: The number of preference pairs that we construct from the 3 datasets in our experiments.
3.1 Dataset Sampling
We utilize a combination of open source instruction models, the GPT-3.5 (Brown et al., 2020) and GPT-4 (OpenAI et al., 2024) series to generate data, ensuring diversity and inclusion of both high-quality and relatively low-quality responses. The open-source models consist of various sizes of the instruction-tuned versions of Llama-3, Llama-3.1, Llama-3.2 (Grattafiori et al., 2024), Llama-2 (Touvron et al., 2023), Qwen-2 (Yang et al., 2024a), InternLM-2 (Cai et al., 2024), and Mistral (Jiang et al., 2023).
In total, we include 12 candidate models for generation. For each question and its corresponding reference in the dataset, we randomly select two models’ generations to form preference pair.
3.2 Dataset Labeling
Dataset Data Example
WebGLM (Liu et al., 2023) Question: Why are different tiers (regular < mid < premium) of gas’ prices almost always 10 cents different? References: [The gap between premium and regular gas has…, According to national averages, the price…] Answer: The 10 cent difference between the different tiers of gas prices is likely due to a convention…
Our Prompt: Answer the following question: {question} Your response should be based on the following passages: {passages} When you respond, you should refer to the source of information…
Yelp (Yelp, 2021) [ Name: The Green Pheasant Address: 215 1st Ave S City: Nashville State: TN Attributes:{ HappyHour: True, DogsAllowed: False, … } ]
Our Prompt: Write an overview about the following business based only on the provided structured data in the JSON format…
XSum (Narayan et al., 2018) Document: The full cost of damage in Newton Stewart, one of the areas worst affected, is still being assessed. Repair work is ongoing in Hawick and many roads… Summary: Clean-up operations are continuing across the Scottish Borders and Dumfries and Galloway after…
Our Prompt: Summarize the following document: {document}…
Table 2: Illustration and statistics of the original datasets and the prompts used to construct the preference data. For WebGLM, LLMs will generate responses based on the reference. For Yelp, LLMs will convert the JSON data into a descriptive overview. For XSum, LLMs will summarize the given document.
We use o3-mini (OpenAI, 2025) to label the data. An illustration of our labeling methods is shown in Figure 2. Given a question and a pair of responses from different models, we prompt o3-mini to compare and select the preferred response.
Specifically, we ask o3-mini to compare the responses based on the four metrics outlined earlier, assessing them individually. In the prompt, we explicitly ask o3-mini to put heavier weights on hallucination and comprehensiveness metrics, as they are crucial to the answer quality, while the other two mainly improves the readability. After the o3-mini has made the individual judgments on the 4 metrics, it will generate an overall preference for the pair data based on the judgments above. We may end up getting a preferred answer with no hallucination but a bit verbose. And this approach acknowledges real-world scenarios where responses are rarely perfect, and trade-offs are often necessary.
4 Dataset Evaluation
4.1 Self Evaluation
WebGLM Yelp XSum Avg
Consist. 97.9 98.8 95.2 97.3
Table 3: The consistency rate for the self-evaluation of the constructed dataset. We measure whether the o3-mini agrees with the comparison results previously made.
Figure 3: An illustration of our self-evaluation method. For a sample of the constructed data pairs, we provide o3-mini with the previous judgment, and ask it to re-evaluate the comparison result.
In this subsection, we evaluate the consistency of the evaluations provided by o3-mini, a key metric commonly used to demonstrate the reliability of the annotations. To objectively assess the quality of our constructed dataset, we design a self-evaluation method that measures the consistency of o3-mini’s responses. An illustration of this method is shown in Figure 3. Specifically, for each prompt, a chosen response, and a rejected response in the dataset, we prompt o3-mini to revisit its previous comparison result and verify whether it maintains its original judgment. We define the consistency rate as the proportion of samples where the evaluation results remain unchanged across both assessments. In this experiment, we randomly select 1,000 samples for re-evaluation.
The results of this experiment are presented in Table 3. We observe a very high consistency rate across the three tasks and the overall consistency rate exceeds 97%. The results demonstrate that o3-mini produces stable and consistent labels for most of the data according to well-defined criteria, and reflects the quality of the dataset.
WebGLM Yelp XSum Avg
Consist. 0.79 0.80 0.83 0.81
Table 4: The consistency rate between human evaluation and the o3-mini labeled dataset.
4.2 Human Evaluation
Models WebGLM Yelp XSum Average
UltraRM-13b (Cui et al., 2024) 71.0 77.4 79.2 75.9
llama-3-tulu-2-8b-uf-mean-rm (Ivison et al., 2024) 71.0 76.2 78.8 75.3
internlm2-7b-reward (Cai et al., 2024) 72.0 73.0 80.4 75.1
Eurus-RM-7b (Yuan et al., 2024) 71.8 74.4 77.8 74.7
FsfairX-LLaMA3-RM-v0.1 (Dong et al., 2024) 71.6 72.4 77.8 73.9
Llama-3-OffsetBias-RM-8B (Park et al., 2024) 70.0 70.4 76.6 72.3
URM-LLaMa-3.1-8B (Lou et al., 2024) 66.8 70.2 76.0 71.0
QRM-Llama3.1-8B-v2 (Dorka, 2024) 68.4 68.0 74.6 70.3
GRM-Llama3.2-3B-rewardmodel-ft (Yang et al., 2024b) 64.6 71.6 73.4 69.9
Skywork-Reward-Llama-3.1-8B-v0.2 (Liu et al., 2024) 64.8 68.6 72.8 68.7
Table 5: The evaluation results of the existing reward models on the 3 tasks. They achieve SOTA performance on chatting, safety, and reasoning evaluation, but do not excel in RAG tasks.
We also performed human evaluations to assess the alignment of AI annotations with human preferences. Specifically, we randomly select 100 samples with paired responses from each dataset and ask the annotators to evaluate using the same pipeline described in Section 3.2 and illustrated in Figure 2. Annotators compare the responses based on each metric and determine the preference pair. We calculate the agreement ratio between the human annotators and o3-mini in the preference pairs, and the results are shown in Table 4. We observe an overall agreement rate of 81%, with consistent agreement across the three tasks. This metric is comparable to the figures reported in Jin et al. (2024), where RAG is evaluated based on helpfulness and harmlessness criteria. These results highlight proprietary LLMs’ ability to effectively capture human preferences in assessing RAG response quality.
5 Limitation of Existing Reward Models
In this section, we evaluate several existing reward models in our test set. We select models from RewardBench (Lambert et al., 2024) known for their strong performance in assessing aspects such as helpfulness, safety, and reasoning. We examine their performance on diverse RAG scenarios using our curated test set (Table 1), and demonstrate their limitation of the evaluation on RAG domains.
The experiment results are shown in Table 5. While many of the listed reward models achieve accuracy higher than 90% in evaluating chat, safety, and reasoning tasks, their overall accuracy in RAG scenarios is below 80%. This underscores a significant gap between mainstream reward models and the unique requirements of RAG tasks. We also observe inconsistent performance across tasks for the reward models, with better results on Data-to-Text and Summarization tasks (e.g. the Yelp dataset and the XSum dataset) compared to the QA task (WebGLM dataset), suggesting that current reward models are not uniformly capable across different RAG scenarios. Interestingly, several reward models that achieve state-of-the-art (SOTA) performance in reasoning and safety evaluations, as shown on the leaderboard [^2] https://huggingface.co/spaces/allenai/reward-bench, do not perform well on RAG tasks. For example, models like URM-LLaMa-3.1-8B (Lou et al., 2024) and Skywork-Reward-Llama-3.1-8B-v0.2 (Liu et al., 2024) underperform in the RAG domain.
In contrast, UltraRM-13b (Cui et al., 2024), which performs suboptimally in reasoning and safety evaluations but excels in assessing helpfulness and instruction-following, achieves the top accuracy on RAG tasks. This suggests that reward models trained primarily on reasoning tasks may not generalize effectively to evaluating RAG-specific generations. Most of the existing reward models could not excel in expressing the preference in RAG scenarios. Domain-specific training data are therefore essential to address this gap and improve RAG performance evaluation.
6 Experiments
We conduct both reward model training and reinforcement learning using our RAG-Reward dataset. In total, 35K preference pairs are used for reward modeling (see Table 1). Additionally, we create a 3K-sample development set for sampling and learning during RLHF training. To evaluate the performance of the policy and reward models, a held-out test set of 1.5K samples is used.
WebGLM Yelp XSum Average
Acc. 81.4 87.6 84.0 84.3
Table 6: The evaluation results of the reward model on the 3 tasks. The accuracy is calculated as the proportion of test samples where the reward model assigns a higher score to the chosen response than to the rejected response.
6.1 Reward Modeling
We adopt the common approach to train the Bradley-Terry reward model (Bradley and Terry, 1952; Ouyang et al., 2022) to learn the reward signal from the preference data. Specifically, we use Llama-3.1-8B-Instruct (Grattafiori et al., 2024) as the base model for training. We train the reward model with a learning rate of $2e^{-6}$, a global batch size of 64, a max length of 4096, and an epoch of 1 on 4 H100-80G GPUs.
During the test stage, each test sample contains a chosen response and a rejected response. The accuracy is calculated as the proportion of test samples in which the reward model assigns a higher score to the chosen response than to the rejected one. The detailed results are shown in Table 6. We observe a high accuracy of 84.3% for the reward model, demonstrating its effectiveness in aligning with the intended criteria. Compared to the strong reward models in Table 5, our model achieves the highest accuracy, highlighting the advantage of leveraging RAG-specific reward data.
Furthermore, we observe a consistent accuracy across the 3 tasks, indicating that the reward model could jointly learn the preference signal from diverse tasks and domains. Notably, the reward model achieves the highest accuracy on the Data-to-Text task, while its performance is relatively lower on the Question-Answering task. This difference suggests that comparing structured data with text data is easier for the reward model, while evaluating the quality of a long-form QA poses a greater challenge. This observation aligns with our intuition and expectations.
6.2 Preference Alignment
Mistral-7B-v0.1 Llama-3.2-3B
WebGLM 64.2 62.6
Yelp 77.8 75.0
Xsum 65.6 64.0
Average 69.2 67.2
Table 7: The RLHF results on the 3 tasks. The win rate is calculated as the proportion of test samples where the reward model assigns a higher score to the response generated by the post-trained policy model.
Mistral-7B-v0.1 Llama-3.2-3B
WebGLM 61.2 60.0
Yelp 80.2 73.6
XSum 63.8 63.2
Average 68.4 65.6
Table 8: The RLHF results on the 3 tasks. The win rate is calculated as the proportion of test samples where the o3-mini prefers the response generated by the post-trained policy model.
We adopt the RAFT algorithm (Dong et al., 2023) to perform the preference alignment. RAFT utilizes the reward model to select the response with the highest reward score from $N$ candidate responses, and then fine-tunes the policy model on this selected set of responses.
We set $N=16$ in our experiments. We use Llama-3.2-3B-Instruct (Grattafiori et al., 2024) and Mistral-7B-Instruct-v0.1 (Jiang et al., 2023) as initial policy models for sampling, followed by RAFT training. Both Mistral and Llama models are fine-tuned with a learning rate of $5e^{-6}$, an epoch of 1, a packing length of 4096, and a global batch size of 16, using the axolotl package [^3] https://github.com/axolotl-ai-cloud/axolotl.
To measure the improvement brought by alignment training, we first sample responses from our held-out test set using both the initial models and the post-trained policy models. Thus, for each prompt, we have paired responses from the two models. These pairs are then evaluated using the reward model, and we calculate the proportion of cases where the responses from the post-trained models are preferred. Furthermore, we ask o3-mini to compare the pairs based on the criteria introduced in Section 3. For both metrics, a baseline of 50% indicates no improvement in the policy models. The calculated scores are referred to as the win rate against the initial model.
The experiment results are shown in Table 7 and Table 8. We observe a clear improvement in the policy models after a single iteration of RAFT. Both the reward model and o3-mini agree that generations align more closely with the RAG metrics, as the average win rate is significantly above 50%. These results highlight the effectiveness of our dataset and the reward model. The ratings across the 3 tasks from the reward model is very similar to the ratings from o3-mini, showing that our reward model learns the rationale of rating from o3-mini.
However, we also observe some imbalance in learning for the policy model across tasks. As shown in the tables, there are differences in the win rate across 3 tasks differences. Specifically, the win rate for Yelp could reach 80% while the other 2 are only above 60%, even though they are trained on the same number of samples for each task. The comparison reveals that the difficulties are different across RAG scenarios.
WebGLM Yelp XSum Average
Human 62.0 70.0 66.0 66.0
Table 9: The human evaluation results of the policy model (Mistral) after RAFT on 3 tasks. The agreement is calculated as the proportion of test samples where the generation after RAFT is preferred by humans.
6.3 Human Evaluation
To further validate the improvement of the policy model, we leverage human evaluation of the generations from it. Specifically, we select the model which trained from Mistral-7B-Instruct with one iteration of RAFT, and the Mistral itself as the reference. We adopt the same evaluation strategy introduced in Section 6.2 and replace the reward model and o3-mini with human annotators. Due to the expense, we select 50 samples of each dataset for human labeling. From Table 9, we observe an agreement far above 50%, indicating the effiveness of the RAFT training to improve the policy model on RAG domains. Compared with the results from the reward model and o3-mini, we discover the same trend across 3 tasks, which further shows the alignment of our dataset with humans.
7 Conclusion
In this paper, we introduce RAG-Reward, a high-quality preference dataset designed for Retrieval-Augmented Generation (RAG). Our dataset is generated through a novel automated AI annotation pipeline, leveraging both open-source and proprietary models to enhance generalization and versatility. To ensure fair and reliable evaluations, we use o3-mini to assess generation quality based on four key metrics carefully selected by human experts. The dataset spans multiple domains, including Question Answering, Data-to-Text, and Summarization, resulting in a large-scale and diverse benchmark. The experimental results show strong alignment with human evaluations, demonstrating the effectiveness of RAG-Reward in reward modeling and reinforcement learning. These findings highlight the potential of our dataset to advance both the evaluation and generation of RAG systems. To foster further research, we will publicly release the dataset to the community.
8 Limitations
In this paper, we constructed a large-scale, high-quality dataset tailored for RAG scenarios and demonstrated the effectiveness of our proposed pipeline through RLHF experiments. However, due to computational constraints, we did not conduct large-scale RLHF training or implement more complex algorithms such as PPO. Future work could explore training larger reward models and incorporating iterative-DPO or PPO to further enhance performance in RAG domains.